OpenStudy (anonymous):

Please help, algebra 2 question Your friend hands you a graph of the performance statistics of the latest powerboat to be produced. He says, “I know this graph is f(x) = –2(x – 4)^3 – 3 but I can’t remember how it is related to the graph of x^3.” Explain to your friend how the graph f(x) is a translation of the graph x^3.

OpenStudy (lena772):

http://www.purplemath.com/modules/fcntrans.htm

OpenStudy (muzzack):

thx @Lena772

OpenStudy (anonymous):

Thank you! @Lena772

OpenStudy (mathmale):

Suppose you start with f(x)=x^3. Multiplying this by 2 will do what to the graph of f(x)=x^3? suppose you begin with f(x)=x^3 and graph this. Then you throw out x and replace it with (x-4) to obtain g(x)=(x-4)^3. How will the graph of g(x) be different from that of f(x0=x^3?

OpenStudy (mathmale):

Now suppose that you begin with either f(x) or g(x), along with their graphs. What happens to the corresponding graph if you create function h(x) by subtracting 3 from either f(x) or g(x)? Once you've answered all three of these questions, see whether you can describe the graph of the function given you by your friend.

OpenStudy (anonymous):

it moves farther down the x axis

OpenStudy (anonymous):

Thank you sooo much, that helps a ton! @mathmale

OpenStudy (mathmale):

My pleasure!